OpenStudy (anonymous):

A gardener is planning to make a rectangular garden with an area of 80ft^2. She has 12 yards of fencing to put around the perimeter of the garden. What should be the dimension of the garden be?

OpenStudy (anonymous):

faye ?

OpenStudy (anonymous):

let the to sides be x nd y, so arey=x*y=80 and x+y=12/2=6 yards, now convert it into feets= 18 feets so x=18-y puttting in 1st equation, y(18-y)=80 clearly the asnwer is 8 nd 10

OpenStudy (anonymous):

share share :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so the dimension is equal to 8 and 10 ?

OpenStudy (anonymous):

yes, in feets

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

dapuk naintindihan nyo kayu na

OpenStudy (anonymous):

faye281

OpenStudy (anonymous):

thank you po dg123